OpenStudy (anonymous):

Find the standard form of -x^2-4x+3

OpenStudy (anonymous):

I can do positive one but not negative ones it's my first... do I just multiply by -1? no right?

OpenStudy (accessdenied):

What is the standard form of a quadratic?

OpenStudy (accessdenied):

i do believe it is "Ax^2 + Bx + C"

OpenStudy (anonymous):

No, it's a(x-h)^2+k

OpenStudy (anonymous):

What you said is general form... see from standard form you can easily find the vertex line of symmetry and minimum or maximum value

OpenStudy (accessdenied):

ahh, okay. i know that form as vertex form. :o

OpenStudy (anonymous):

Oh, where are you from... lol nvm doesn't matter I really need help with this >.<

OpenStudy (accessdenied):

in the US... Okay, so y = -x^2 - 4x + 3 We'd factor out a -1 from the first two terms y = -(x^2 + 4x) + 3 We want to complete the square for -( (x^2 + 4x + __) - ___ )

OpenStudy (accessdenied):

to complete the square, we'd just divide the 4 coefficient on the x by 2 and then square the value. y = -( (x^2 + 4x + (4/2)^2) - (4/2)^2 ) + 3 = -((x^2 + 4x + 4 - 4) + 3 = -((x+2)^2 - 4) + 3 distribute the -1 back in = -(x+2)^2 + 4 + 3 = -(x+2)^2 + 7

OpenStudy (anonymous):

I'm an idiot!!! Why didn't I just take the negative out! Thank you very much man you saved my life I got kicked out of my schools basketball team because I was messing up... time to come back!

OpenStudy (accessdenied):

no problem! and good luck. :D most things in algebra have the same process needed to figure similar problems out, i think